OpenStudy (liv1234):

Can someone check my answer please? The one I have clicked is my answer choice.

OpenStudy (liv1234):

@welshfella

OpenStudy (liv1234):

@pooja195

OpenStudy (welshfella):

sum of 15 terms = 2 * 1 - 0.5^15 --------- 1 - 0.5

OpenStudy (luckyness):

I believe its correct, Although I would check with Welsh :)

OpenStudy (liv1234):

I got 1.5

OpenStudy (liv1234):

@welshfella

OpenStudy (welshfella):

No its not 1.5 use the formula work out 1 - 0.5^15 divide this by 0.5 then multiply by the first term 2 use your calculator it might not give you the exact answer - but very close

OpenStudy (welshfella):

the formula is S = a1 ( 1 - r^n) ------- 1 - r in your case a1 = 2 , n = 15 and r = 1/2 (or 0.5)

OpenStudy (liv1234):

Okay, but I keep getting 1.5

OpenStudy (welshfella):

what tdo you get for 1 - 0.5^n ?

OpenStudy (welshfella):

- you work out the parentheses first

OpenStudy (welshfella):

i used my calculator and it give me 0,.999694824 but we can call that 1 so multiply 1 by 2 then divide by 0.5

OpenStudy (welshfella):

Sum = 2 * 1 ----- 0.5

OpenStudy (welshfella):

2 / 0.5 = ?

OpenStudy (welshfella):

just work that out on your calculator id you are not ccconfident to do it manually.

OpenStudy (liv1234):

I got 4

OpenStudy (liv1234):

Sorry for the late response, OpenStudy was being slow

OpenStudy (welshfella):

thats right thats the naswer

OpenStudy (liv1234):

Thank you!

OpenStudy (welshfella):

naswer - lol!!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

its worth trying to remember that formula for sum of n terms of a Geometric series.